Grade 12Mechanics

once ignited, a small rocket motor on a spacecrsft exerts a constant force of 10 N for 7.80 seconds. during the burn the rocket causes the 100 kg craft to accelerate uniformly. determine that acceleration.

madona banhan
8 Years agoGrade 12

2 Answers

Arun
8 Years ago
Dear student
 
Force = mass * acceleration 
So acceleration = Force / mass = 10 / 100 m/s^2 = 0.1 m/s^2
 
Regards
Arun (askIITians forum expert)
Khimraj
8 Years ago
Acceleration is given by a = F/m.
F = 10N
m = 100kg
a = 10/100  = 0.1m/s2
Hope it clears.